can you please help me with {x+y=12
{x-y=2

Found 2 solutions by MathGrace, pwac:
Answer by MathGrace(6)

There are several ways to do this problem (linear combination, substitution, graphing). The easiest is by linear combination. You are going to subtract the two equations.
x+y=12
-(x-y=2) Put the common terms over each other like a subtraction problem.
---------
x+y=12
-x+y=-2 Distribute the negative to both sides of equation
---------
0x+2y=10 Combine like terms.
2y=10
2y/2=10/2
y=5

Answer by pwac(253)

Add the two equations to give:
2x=14
x=7 substitute this into 1st equation
7+y=12 subtract 7 from both sides
y=5
Pete